How would you go about solving this factor?

Solve: 8x^2 + 2x = 3

Answer by stanbon(48568)

Solve: 8x^2 + 2x = 3
8x^2+2x-3=0
8x^2+6x-4x-3=0
2x(4x+3)-(4x+3)=0
(4x+3)(2x-1)=0
x = -3/4 or x = 1/2
=================
Cheers,
Stan H.